**TITLE** ASP Conference Series, Vol. **VOLUME**, **PUBLICATION YEAR** **EDITORS**

Publications of the Astronomical Society of the Pacific Conference Series—Instructions for Authors and Editors Using L^{A}TEX Markup

Iman Author

The Name of My Institution, The Full Address of My Institution, Run in Address to Make Minimum Number of Lines

Ima Co-Author

The Name of My Institution, The Full Address of My Institution

Abstract. A concise abstract is recommended. Enter the text of the abstract in an “abstract” environment, i.e., within \begin{abstract} and \end{abstract} commands. Do not make a paragraph break between \begin{abstract} and the first line of the text of the abstract! Abstracts are required for all papers.

1. Introduction

In order to ensure that papers received for publication from different authors are consistent in format, style, and quality, authors are required to type their manuscripts exactly according to the following instructions. The editors will modify the electronic manuscripts as necessary to ensure that they conform to these standards in order to produce a final camera ready copy.

This document is a revised version of the PASP documents paspman.tex and revpaspman.tex^{1}.

2. L^{A}TEX Markup Commands

Authors using L^{A}TEX are requested to use the newpasp style file, and declare it as a substyle to the standard L^{A}TEX article style. A copy of the style file and the file for these instructions are available via anonymous ftp at physics.byu.edu. You will find the files in the subdirectory /pub/pasp/style/new/aspconf.

You should use only those markup commands from L^{A}TEX plus the several extensions provided by this style file. Do not define any commands of your own for any reason (no \def or \newcommand statements).

1The present version, based on L^{A}TEX 2.09, will shortly be revised to incorporate L^{A}TEX 2_{ε}.

2.1. Preamble

The first piece of markup in the manuscript must declare the overall style of the document.

     \documentstyle[11pt,newpasp,twoside]{article}

The \documentstyle command must appear first in any L^{A}TEX file, and this one specifies the main style to be the L^{A}TEX article style using eleven point fonts, with modifications and additions for the newpasp substyle.

To get the running headers at the top of each page (authors’ names on the left page and title on the right page), enter the following commands directly beneath the \documentstyle command:

\markboth{author’s surnames}{title of contribution}

     \pagestyle{myheadings}

The author must include a

     \begin{document}

command to identify the beginning of the main portion of the manuscript.

2.2. Title, Byline, Abstract, etc.

Title and author identification are by way of the standard L^{A}TEX commands \title and \author. An author’s principal affiliation is specified with a separate macro \affil. Each \author command should be followed by a corresponding \affil (address).

\title{lucid text}

     \author{name(s)}
     \affil{address}

The \affil command should be used to give the author’s full postal address. The address will be broken over several lines automatically; do not use L^{A}TEX’s \\ command to indicate the line breaks. Please use mixed case text for all these fields rather than supplying all capitals; the style file will convert to upper case as necessary.

The article’s abstract should be enclosed in an abstract environment.

     \begin{abstract}

abstract text

     \end{abstract}

Don’t include the word “Abstract” in your text; it is inserted automatically. Do not make a paragraph break between \begin{abstract} and the first line of the text of the abstract!

2.3. Sections

The L^{A}TEX article environment supports three levels of sectioning.

     \section{heading}
     \subsection{heading} \subsubsection{heading}

Please use mixed case text for the section heads. Note that these commands delimit sections by marking the beginning of each section; there are not separate commands to identify the ends.

If one wishes to have an acknowledgments section, it should be set off simply with the command

     \acknowledgments

2.4. Equations

Displayed equations can be typeset in many ways using the standard displayed math environments of L^{A}TEX; these three are probably of greatest use.

     \begin{displaymath}
     \end{displaymath}
     \begin{equation}
     \end{equation}
     \begin{eqnarray}
     \end{eqnarray}

The displaymath environment will break out a single, unnumbered formula. The equation will appear the same if it is set in an equation environment, and it will be autonumbered by L^{A}TEX. In order to set several formulæ in which vertical alignment is required, use the eqnarray environment.

2.5. Tables

Tables should appear in table environments.

     \begin{table}
     \caption{text}
     \begin{tabular}{cols}
     \end{tabular}
     \end{table}

There should be only one table per environment. The table environment encloses not only the tabular material but also any title (caption) or footnote information associated with the table. Tabular information is typeset within L^{A}TEX’s tabular environment; the cols argument specifies the formatting for each column. Tables and figures will be identified with arabic numerals, e.g., “Table 2.”; the identifying text, including the number, is generated automatically by the \caption command.

There is a \tableline command for use in tabular environments.

     \tableline

This command produces a single horizontal rule. There should be a \tableline above and below between the column headings, and two at the end of the table. Authors should not use additional \tablelines themselves, and are discouraged from using vertical rules unless essential.

2.6. EPS Files

Authors who can prepare computer graphics in Encapsulated PostScript (EPS) format may use one of two additional markup commands to mark the point of inclusion, both of which should be used inside a L^{A}TEX figure environment. If the DVI translator dvips by Tom Rokicki is available on your computer, it is also possible to prepare the final copy with such figures in place. If you have Tom’s dvips program, you must include the epsf substyle in the \documentstyle definition:

     \documentstyle[11pt,newpasp,twoside,epsf]{article}

You may have to copy epsf.tex to epsf.sty or create a symbolic link, since dvips is distributed with epsf.tex which serves both plain TEX and L^{A}TEX. The same file works fine as a L^{A}TEX style file, it just needs the .sty extension.

The markup commands for including EPS files are:

     \plotone{file}
     \plottwo{file}{file}

The file argument is used to name the file(s) to be included. The \plotone command includes one figure that is scaled to the width of the current text column; \plottwo inserts two figures side by side, and the pair is scaled to fit the text width. If one uses these macros, the necessary vertical space is provided automatically.

     \begin{figure}

\plotone{mygraph.eps} \caption{My EPS graphic.}

     \end{figure}

or

     \begin{figure}

\plottwo{mygraph.eps}{another.eps} \caption{Two related graphics.}

     \end{figure}

Please note that the caption will be centered under the pair of graphics when \plottwo is used. It is not possible to caption the two plots individually with this package at this time. As with tables, figures will be identified with arabic numerals, e.g., “Figure 1.”.

If you must fuss with the positioning and scaling of the EPS plot on the printed page, you can try using this command:

\plotfiddle{file}{vsize}{rot}{hsf }{vsf }{htrans}{vtrans} vsize vertical white space to allow for plot, any valid

L^{A}TEX dimension

rot  rotation angle, in degrees, counter-clockwise
hsf  horiz scale factor, percent
vsf  vert scale factor, percent
htrans horiz translation, in PS points 72/in
vtrans vert translation, in PS points 72/in

If you can produce EPS but you do not have dvips, you can still put the \plotone or \plottwo commands in the the appropriate places, but you will have to comment them out and put in a \vspace{dimen} command to open up the text. The dvips program is in the public domain and is available from labrea.stanford.edu.

A special note to authors: Color EPS files should be avoided if possible. And since it is sometimes necessary to edit EPS files to make them printable, authors should try to avoid EPS files with lines longer than 1024 characters.

2.7. Pasted in Illustrations

Illustrations must be inserted in the text at the appropriate places, with the relevant caption underneath each. The finished pages are reduced by 10% before printing. Thus, illustrations will appear somewhat smaller in print.

These illustrations should appear in figure environments.

     \begin{figure}
     \vspace{dimen}
     \caption{text}
     \end{figure}

There should be only one figure per environment. Space for the figure is created with the \vspace command; dimen should be a valid L^{A}TEX dimension, e.g.,

“2.5in”.

Stick glue is normally preferred for pasting in illustrations, but other adhesives will suffice. Do not use a glue which will wrinkle the paper. Illustrations must not be fixed to the pages by transparent adhesive (cellophane) tape since this will show in the printing. Paste in each illustration so that it is centered on the page. Also check (e.g., with a set-square) that the illustration is not crooked.

Size of Illustrations The maximum width of an illustration is normally 13.4cm (5.25in) so that it will fit within the width of the text area. Of course an illustration may be smaller if appropriate. A large illustration may be placed sideward (“landscape”) on the paper if necessary.

Halftone Illustrations (Photographs) Good glossy original prints are required, black and white only; color plates cannot be reproduced. Photographs cut from other publications will not reproduce well, and usually infringe copyright. Supply both a large photograph and a reduced version pasted in the text. The pasted-in version will be used for size and placement only; a copy machine can be used for this reduction. If a reduction is not supplied, leave appropriate space in text above the figure caption. The publisher will take care of the photographic reduction and mounting of the original glossy print in the space provided by your pasted-in version.

Line Illustrations If line drawings are submitted as the original, they should be drawn in black ink on good quality tracing paper or a glossy photographic print.

The lettering on line illustrations must be satisfactory — not too large, not too small. The preferred style of lettering is “all capitals”, with capital letters between 2mm and 3mm in height (roughly, 6pt to 8pt). Lettering should be done with a lettering stencil or press-on lettering; free-hand or typewritten lettering is not recommended. All lettering and fine detail on both line and photographic illustrations must be very clear and of good visual quality. It is often easier to prepare line illustrations at a larger size (e.g., twice the required size), then have them reduced photographically for pasting into the text.

Illustrations produced on laser printers are acceptable and are probably the preferred way to produce drawings by most authors.

3. References

3.1. In the Text

The reference system to be followed is the standard ApJ system. We adopt the editorial convention of not italicizing the phrase “et al.”.

Single Author Author name followed by the year in parentheses, as in Abt (1990), or author and year both in parentheses (Abt 1990).

Two Authors Author names separated by an ampersand (no comma). Not to be abbreviated subsequently to “Author1 et al.” In parentheses use “(Author1 & Author2 1999)”.

Three Authors When first mentioned in text use the format “Author1, Author2, & Author3 (year)” (note comma before ampersand). Subsequently, abbreviate to “Author1 et al.”. When authors and year are both within parentheses use “(Author1, Author2, & Author3 1999)” or “(Author1 et al. 1999)”.

More than Three Authors Use the format “Author1 et al.”. Please do not use “Author1, Author2, Author3, et al.”, since Authors 2 and 3 will be deleted during the copy editing! In parentheses: “(Author1 et al. 1999)”.

Citing Multiple Works Inside parentheses, citations are separated by a semicolon: “(Biretta, Lo, & Young 1982; Forrest et al. 1987; Lee 1995)”.

Outside of parentheses, use commas: “According to Biretta, Lo, & Young (1982), Forrest et al. (1987), and Lee (1995), there is strong evidence to suggest that...”.

3.2. Reference List

There is a references environment that sets off the list of references and adjusts spacing parameters.

     \begin{references}

\reference bibliographic information

        .
        .
     \end{references}

The bibliographic information should be in the order directed by Abt (1990): author, year, journal, volume, and page. For instance, the reference for this editorial would be typed in as

     Abt, H. 1990, ApJ, 357, 1

Note that there is not a comma following the author name(s), there is not a trailing period at the end of the reference, and the entire line is set in the body typeface (no font changes).

Please note the order of the bibliographical information in the following entries:

     Biretta, J. A., Lo, K., & Young, P. J. 1982 in AIP Conf.
     Proc., 83, The Galactic Center, ed. G. R. Riegler & R. D.
     Blandford (New York: AIP), 91
     van der Kruit, P. C., & Shostak, G. S. 1983, in IAU Symp.
     100, Internal Kinematics and Dynamics of Galaxies, ed. E.
     Athanassoula (Dordrecht: Reidel), 69
     Garcia-Lorenzo, B., Mediavilla, E., Arribas, S., & del
     Burgo, C. 1998, in ASP Conf. Ser. Vol. 152, Fiber Optics
     in Astronomy III, ed. S. Arribas, E. Mediavilla \& F.
     Watson (San Francisco: ASP), 185

Care should be taken that each literature citation in the manuscript has its counterpart in the reference list and vice versa. Care should also be given to checking the accuracy of the references—author(s), date, volume, and page number. While the accuracy of the references is the sole responsibility of the author(s), the editor(s) should nevertheless aim to ensure uniformity of presentation throughout the volume.

3.3. Abbreviations for Journals

There are macros for many of the oft-referenced journals so that authors may use the L^{A}TEX names rather than having to look up a particular journal’s specific abbreviation. Any stylistic requirements of the editors are taken care of by the macros, so authors need not be concerned about such editorial preferences.

8 Author & Co-author

\aj  Astronomical Journal (AJ)
\araa  Annual Review of Astronomy and Astrophysics(ARA&A)
\apj  Astrophysical Journal (ApJ)
\apjs  , Supplement Series (ApJS)
\ao  Applied Optics (Appl.Optics)
\apss  Astrophysics and Space Science(Ap&SS)
\aap  Astronomy and Astrophysics (A&A)
\aaps  , Supplement Series (A&AS)
\azh  Astronomicheskii Zhurnal (AZh)
\baas  Bulletin of the AAS (BAAS)
\jrasc Journal of the RAS of Canada (JRASC)
\memras Memoirs of the RAS (MmRAS)
\mnras Monthly Notices of the RAS (MNRAS)
\pra  Physical Review A: General Physics (Phys.Rev.A)
\prb  Physical Review B: Solid State (Phys.Rev.B)
\prc  Physical Review C (Phys.Rev.C)
\prd  Physical Review D (Phys.Rev.D)
\prl  Physical Review Letters (Phys.Rev.Lett)
\pasp  Publications of the ASP (PASP)
\pasj  Publications of the ASJ (PASJ)
\qjras Quarterly Journal of the RAS (QJRAS)
\skytel Sky and Telescope (S&T)
\sovast Soviet Astronomy (Soviet Ast.)
\ssr  Space Science Reviews (Space Sci.Rev.)
\zap  Zeitschrift für Astrophysik (ZAp)

3.4. Mathematical Symbols The following macros produce various troublesome or laborious mathematical symbols.

          \deg  (^{∘}) \farcs  (.′′)
          \sun  (⊙) \fp  (.^{p})
          \earth (⊕) \micron  (μm)
          \la  (∼ <) \onehalf  (^{1}/_{2})
          \ga  (∼ >) \onethird  (^{1}/_{3})
          \arcmin (′) \twothirds  (^{2}/_{3})
          \arcsec (′′) \onequarter  (^{1}/_{4})
          \fd  (.^{d}) &\threequarters (^{3}/_{4})

\fh  (.^{h}) \ubvr  (UBVR)
\fm  (.^{m}) \ub  (U −B)
\fs  (.^{s}) \bv  (B−V)
\fdg  (.^{∘}) \vr  (V −R)

          \farcm (.′) \ur  (U −R)

4. Instructions to Editors

Please observe the following guidelines.

All contributions must carry the eyebrow slug (giving the title of the conference, the ASP Conf. Ser. volume number and the name(s) of the editor(s). The slug will appear in the top lefthand corner of the first page of the contribution. To achieve this, open the file newpasp.sty and search for the lines

\def\@jourvol{**VOLUME**}
\def\cpr@year{**PUBLICATION YEAR**}
\def\vol@title{**TITLE**}
\def\vol@author{**EDITORS**}

and replace “**VOLUME**”, etc., by the relevant information.

Remember to check the volume number and publication date with Enid Livingstone at PASP@io.ln.byu.edu before printing out the final copy!

Check that the running headers are inserted and alternate correctly (has the author remembered to put the parameter “twoside” in the documentstyle command?).

Ensure that the pagination for each article is correct. To change the pagination use \setcounter{page}{number}, where number is the first page number of the article. This command should be inserted in each article file after the

\pagestyle{myheadings} command.

Standardize the bibliographical referencing as outline above.

Ensure that the page width is never exceeded. Any text, tables or illustrations that stretch beyond the margins will be truncated in the printing process.

5. Persons to Contact

For details concerning the volume number allocated to your proceedings volume, date of publication, questions relating to the book production process, and administrative details, please contact Enid Livingstone

     (PASP@io.ln.byu.edu).

All queries relating to the L^{A}TEX style file, author’s/editor’s instruction manual, etc., should be directed to Terry Mahoney (tjm@ll.iac.es).